QUESTION 4 Given f(x)=2x-5 and g(x)= x2 + x - 20, find each: a) [fog)(3) b)[gºf](x) IT I Arial 3(12pt » T.:= . 31.25
Answer #1

son . . Consider . Given f(x) = 2x-5, g(x) = x²+x-20 © (fog)(3) , this is composition of two - function . feg . so. [tog] (u)(b) f(x) = 2x-5 f g(x)= x²+2-20 son Now [701](a) = [9127-537 = (27-5)2 + (2x-3) -20 = 4x²+25-20x + 2x-5.-2. = 4x²_18*+/5 - 2x